# Robotic Policy Learning via Human-assisted Action Preference Optimization

### v2

@@ 4.3 Generalization to Novel Tasks
In this section, we assess our framework’s generalization capability under three novel scenarios, as illustrated in Figure 2. (1) Position Disruption: For the Square_D0 task, we replace the fixed initial stick position with randomized placements within a bounded operational area. (2) Background Disruption: In the StackThree_D0 task, we substitute the default white background with a gray one. (3) Texture Disruption: In the ThreePiece_D0 task, the original red blocks are transmuted to wood-grain visual properties. These experiments systematically evaluate robustness against spatial, background, and visual texture variations. To fine-tune the base model on novel disruption scenarios, we collect 20 interaction trajectories under disruption scenarios and combine them with 20 expert demonstrations from the original task for subsequent fine-tuning.
Our objective is to develop a human-assisted action preference optimization method that facilitates continuous improvement, enabling performance enhancements in novel disruption scenarios while retaining original task capabilities during model fine-tuning. Thus, we evaluate the performance of the fine-tuned model across both disruption scenarios and original scenarios.
As shown in Tabel 4.2, the base policy exhibits some degree of performance degradation in disruption scenarios. However, the performance decline is relatively minor in cases of background disruption, whereas disruptions in object texture and position significantly impact performance. Both behavior cloning methods and preference optimization methods struggle to achieve significant performance improvements in novel disruption scenarios. In contrast, our approach can effectively adapt to new disruption scenarios through adaptive reweighting.
Table  4.2 presents the performance of the optimized model after being fine-tuned from disruption data on the original task. The results reveal that behavior cloning methods exhibit severe catastrophic forgetting, resulting in substantial performance degradation. By contrast, the preference optimization method achieves mitigated performance decline with the constraints of the reference model.
Besides, our approach utilizes adaptive reweighting to effectively integrate knowledge from both expert demonstrations and interaction trajectories. This mechanism not only facilitates learning from diverse data sources but also leads to improved performance on the original task.
@@ 4.4 The Performance of Lifelong Learning
To investigate whether our method can iteratively improve via environment interaction, we deploy our model to interact with environments while updating the model every 20 interaction rollouts. Additionally, we provide comparison results using a behavior cloning policy trained with the same number of expert demonstrations as our baseline. For each updated model, we conduct 50 trials and report the success rate.
As shown in Figure 3(a-b), our method achieves superior performance compared to the baseline, demonstrating its ability to effectively leverage sub-optimal human intervention trajectories for iterative model improvement. When the base policy exhibits diminishing improvement with increasing expert demonstrations, our method enables continual performance gains from the interaction trajectories. Besides, this improvement trend is accompanied by a corresponding reduction in the required human intervention ratio, as shown in Figure 3(c)
Figure 3: Lifelong learning results of our mehtod.
@@ 4.5 Generalization to various VLA models
To validate that our method can be adapted to different VLA models, we applied our approach to fine-tune the $\pi 0$-FAST [31] model. $\pi 0$-FAST applies discrete cosine transform encoding to encode the action chunking into discrete tokens for VLA training. To adopt this model for downstream tasks, We regenerate the action tokenizer with 5 action chunking step for each task.
As shown in Table 9, the base model could achieve a higher success rate compared with the OpenVLA model, benefiting from its ability to predict action chunking for robotic manipulation. Further, we compare our method with both the behavior cloning method and preference optimization method, the results demonstrate that our method could achieve consistent improvement for the $\pi 0$-FAST fine-tuning. The results prove that our method could our method can be applied to the fine-tuning of various VLA models, achieving consistent performance gains.
@@
Methods Coffee_D0 StackThree_D0
Base policy 68% 64%
Dagger 64% 66%
TPO 48% 52%
Ours 76% 74%
@@ 4.6 Real-world Experiments
In this work, we conduct the challenging fine-grained robotic manipulation task “Insert the square into the stick” as shown in Figure 7(a), which requires the robot to grasp the square and precisely insert into the stick. To collect expert demonstrations, we utilize the spacemouse device to gather 100 high-quality trajectories at an action frequency of 20 Hz. We fine-tune the OpenVLA model with the collected demonstration as the base model. Further, we deploy the base model to interact with environments and propose the real-time human-in-the-loop interventions to collect 20 interaction trajectories for subsequent preference alignment optimization. To fairly compare performance, all methods are evaluated under the same experimental setup, and we report the average success rate from 20 trials.
To comprehensively evaluate our method, experiments are conducted not only under in-distribution but also across three distinct disruption settings as shown in Figure 7(b-e): (1) Position Disruption: We change the position distribution of the stick. (2) Background Disruption: We replace the tablecloth from white to brown. (3) Texture Disruption: We replace the black stick to the green one.
As demonstrated in Table 4.5, our method demonstrated robust adaptability to these downstream disruption scenarios. The results empirically validate the method’s practical utility for real-world deployment in unstructured environments.
@@ 5 Conclusion
In this work, we introduce the human-assisted action preference optimization method, consisting of two critical components: a human-robot collaboration framework for reliable deployment and an action preference optimization process with adaptive reweighting for stable VLA model optimization. Through our method, we could promote continuous improvement during the deployment of VLA models. We hope our method could bring insights for efficient and effective VLA model adaptation on downstream manipulation tasks.
Limitation. While our work study the preference alignment optimization for VLA models, the experiments are based solely on autoregressive VLA models. Future work should explore a broader range of VLA frameworks, including regression-based approaches and diffusion policy models, to ensure the generalizability of our method across different architectures.

### v3

@@ 4.3 Generalization to Novel Tasks
In this section, we assess APO’s generalization capability under three novel scenarios, as illustrated in Figure 3. (1) Position Disruption: For the Square_D0 task, we replace the fixed initial stick position with randomized placements within a bounded operational area. (2) Background Disruption: In the StackThree_D0 task, we substitute the default white background with a gray one. (3) Texture Disruption: In the ThreePiece_D0 task, the original red blocks are transmuted to wood-grain visual properties. These experiments systematically evaluate robustness against spatial, background, and visual texture variations. To fine-tune the base model on novel disruption scenarios, we collect 20 interaction trajectories under disruption scenarios and combine them with 20 expert demonstrations from the original task for subsequent fine-tuning.
Our objective is to develop an action preference optimization method that facilitates continuous improvement, enabling performance enhancements in novel disruption scenarios while retaining original task capabilities during model fine-tuning. Thus, we evaluate the performance of the fine-tuned model across both disruption scenarios and original scenarios.
As shown in Tabel 4.2, the base policy exhibits some degree of performance degradation in disruption scenarios. However, the performance decline is relatively minor in cases of background disruption, whereas disruptions in object texture and position significantly impact performance. Both behavior cloning methods and preference optimization methods struggle to achieve significant performance improvements in novel disruption scenarios. In contrast, APO can effectively adapt to new disruption scenarios through adaptive reweighting.
Table 4.2 presents the performance of the optimized model after being fine-tuned from disruption data on the original task. The results reveal that behavior cloning methods exhibit severe catastrophic forgetting, resulting in substantial performance degradation. By contrast, the preference optimization method achieves mitigated performance decline with the constraints of the reference model.
Besides, APO utilizes adaptive reweighting to effectively integrate knowledge from both expert demonstrations and interaction trajectories. This mechanism not only facilitates learning from diverse data sources but also leads to improved performance on the original task.
@@ 4.4 The Performance of Lifelong Learning
To investigate whether APO can iteratively improve via environment interaction, we deploy APO to interact with environments while updating the model every 20 interaction rollouts. e provide comparison results using a behavior cloning policy trained with the same number of expert demonstrations as our baseline. For each updated model, we conduct 50 trials and report the success rate.
As shown in Figure 4(a-b), APO achieves superior performance compared to the baseline, demonstrating its ability to effectively leverage sub-optimal human intervention trajectories for iterative model improvement. When the base policy exhibits diminishing improvement with increasing expert demonstrations, APO enables continual performance gains from the interaction trajectories. Besides, this improvement trend is accompanied by a corresponding reduction in the required human intervention ratio, as shown in Figure 4(c)
@@ 4.5 Generalization to various VLA models
To validate that APO can be adapted to different VLA models, we applied APO to fine-tune the $\pi 0$-FAST [33] model. $\pi 0$-FAST applies discrete cosine transform encoding to encode the action chunking into discrete tokens for VLA training. To adopt this model for downstream tasks, We regenerate the action tokenizer with 5 action chunking step for each task.
As shown in Table 8, the base model could achieve a higher success rate compared with the OpenVLA model, benefiting from its ability to predict action chunking for robotic manipulation. Further, we compare APO with both the behavior cloning method and preference optimization method, the results demonstrate that APO could achieve consistent improvement for the $\pi 0$-FAST fine-tuning. The results prove that APO could be applied to the fine-tuning of various VLA models, achieving consistent performance gains.
@@
Methods Coffee_D0 StackThree_D0 Insert Square
Base policy 68% 64% 85%
Dagger 64% 66% 85%
TPO 48% 52% 90%
APO 76% 74% 95%
@@ 4.6 Real-world Experiments
In this work, we conduct the challenging fine-grained robotic manipulation task “Insert the square into the stick” as shown in Figure 8(a), which requires the robot to grasp the square and precisely insert into the stick. To collect expert demonstrations, we utilize the spacemouse device to gather 100 high-quality trajectories at an action frequency of 20 Hz. We fine-tune the OpenVLA model with the collected demonstration as the base model. Further, we deploy the base model to interact with environments and propose the real-time human-in-the-loop interventions to collect 20 interaction trajectories for subsequent action preference optimization. All methods are evaluated under the same experimental setup, and we report the average success rate from 20 trials. For a comprehensive evaluation in real-world scenarios, APO was also tested on the “hang cup on the rack” and “put lemon on the plate” tasks.
To comprehensively evaluate APO, experiments are conducted not only under in-distribution but also across three distinct disruption settings as shown in Figure 8(b-e): (1) Position Disruption: We change the position distribution of the stick. (2) Background Disruption: We replace the tablecloth from white to brown. (3) Texture Disruption: We replace the black stick to the green one.
As demonstrated in Table 4.5, APO demonstrated robust adaptability to these downstream disruption scenarios. The results empirically validate the method’s practical utility for real-world deployment in unstructured environments. We also adopt APO to fine-tune the $\pi_{0}$-FAST model in the real world scenario. The results in Table 8 prove APO could achieve consistent performance gains over other methods.
4.7 Correction from Failure scenarios
Figure 6: The rollout trajectory of APO. As indicated by the bold red and green boxes, APO can autonomously correct form failure scenarios.
In this work, we propose the APO method that enables models not only to avoid failure modes but also to self-correct within failure scenarios. As shown in Figure 6, we provide examples of failure correction across multiple tasks, demonstrating the corrective strategies learned by APO.
For instance, in Figure 6(a,b), when the model initially fails to grasp an object, APO identifies the failure and initiates a re-grasp attempt. Similarly, in Figure 6(c), when a precise insertion operation is obstructed, APO learns to iteratively adjust its gripper position until the insertion is successfully completed. These examples illustrate that APO has successfully learned to recover from common failure scenarios, which directly contributes to its improved overall performance.
@@ 5 Conclusion
In this work, we introduce the Action Preference Optimization (APO) method to fully exploit valuable information in failure trajectories while maintaining the stability required for large-scale VLA models. This method builds on a human-robot collaboration framework for reliable deployment, and utilizes an adaptive reweighting preference optimization algorithm with action-level binary desirability signals for stable VLA model optimization. Through APO, we could promote continuous improvement during the deployment of VLA models. We hope APO could bring insights for efficient and effective VLA model adaptation on downstream manipulation tasks.
Discussion and Future Work. While our work study the preference alignment optimization for VLA models, the experiments are based solely on autoregressive VLA models. Future work should explore a broader range of VLA frameworks, including regression-based approaches and diffusion policy models, to ensure the generalizability of our method across different architectures.